[Brahms-l] Shift report 20050515 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Sun May 15 2005 - 07:41:14 EDT
Shifter leader was Michael
Shifters were 


<pre>
Good Beam from midnight until 4am, then after dump problem with
Oxygen monitor. 

Took runs 
14654 14659 14660 14661 14662 14663  	
  T1     T2        T3      T4      T5
757969 	107079 	16676 	13715 	137717 	

For these runs I misstyped the D1 current as 3440 so they need to be updated in the database. Also from run 14659 I started it before I had finished 14654 in E log. 

I had one problem with DAQ. Runserver and IOserver got hung but I  
restarted them thanks to procedure in handbook. After some time
I had problem getting data with Supermon. 

================================

Polarization 
Time   Blue      Yellow           
1am  -48.7+-2.8  -44.9+-2.8
3am  -52.3+-2.8  -44.1+-2.7
4am  -49.4+-2.0  -22.9+-2.8
</pre>

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun May 15 07:41:26 2005

This archive was generated by hypermail 2.1.8 : Sun May 15 2005 - 07:41:39 EDT